1. Evaluate D(sin(x²+3x-1))
2. Given that f(x)=3x-1
g(x)=cosx
h(x)=e^2x+1
Find (fogoh)x and (hogof)x
3. Evaluate lim[⅔x] x->6-, [•] in the greatest integer formula.
4. Find lim (x(x+1)(2+x))/x+x³
5. Find the different coefficient of (i) xlog5^x (ii) x³log8^x
1
Expert's answer
2019-09-04T09:44:23-0400

1.

"D(\sin(x^2+3x-1))=(x^2+3x-1)'\cos(x^2+3x-1)dx=(2x+3)\cos(x^2+3x-1)dx"

2.

"f\circ g\circ h(x)=f(g(h(x)))=f(\cos(e^{2x}+1))=3\cos(e^{2x}+1)-1""h\circ g\circ f(x)=h(g(f(x)))=h(\cos(3x-1))=e^{2\cos(3x-1)}+1"

3.

"\lim\limits_{x\to6-}[\frac{2x}{3}]=3" - For all 3<x<6 "\left[\frac{2x}{3}\right]=3" . Here [x] is the greatest integer function.

4.

"\lim\frac{x(x+1)(x+2)}{x+x^3}" no information on x here. If "x\to\infty" then "\lim\limits_{x\to \infty}\frac{x(x+1)(x+2)}{x+x^3}=\lim\limits_{x\to \infty}\frac{x^3+3x^2+2x}{x+x^3}=1"

5.

"\frac{d}{dx}x\log5^x=\log5^x+x\frac{5^x\log5}{5^x}=\log5^x+x\log5"

"\frac{d}{dx}x^3\log8^x=3x^2\log8^x+x^3\frac{8^x\log8}{8^x}=3x^2\log8^x+x^3\log8"